You are provided with the information below with respect to the possible returns of Mariakani Ltd....

  1. You are provided with the information below with respect to the possible returns of Mariakani Ltd.

Probability

0.1

0.3

0.4

0.2

Return (%)

23

18

12

7

You are also given that the expected market return is 10.4%, the standard deviation of the market returns is 2.25%, whereas the correlation between the returns of Mariakani Ltd and the market is 0.84

Required

  1. Determine the expected return and risk for Mariakani Ltd                         (6 marks)
  2. Compute the beta factor for Mariakani Ltd and interpret its meaning          (4 marks)
Answer #1

A B C D E F G H I J J K L M N O P 1 b d=b - 13.9 frate Probability Return - ER Prob*Square 0.10 0.30 0.40 0.20 c =a*b ER = Re